OpenStudy (toxicsugar22):

hi can any1 help me solve this 3^(9-3x)=27

OpenStudy (toxicsugar22):

and give me some promblems related to this so i know how to do it for my test tommorow

OpenStudy (toxicsugar22):

please

OpenStudy (toxicsugar22):

hello r u therer

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

Convert 27 to base 3 and some exponent.

OpenStudy (anonymous):

which is nine

OpenStudy (anonymous):

its taking 9 to the 3 power which is 729

OpenStudy (anonymous):

The base must be 3.

myininaya (myininaya):

27=3^3

myininaya (myininaya):

so 9-3x=3

OpenStudy (anonymous):

9-3x=3

OpenStudy (toxicsugar22):

thanks